OpenStudy (anonymous):

Solve for x in the proportion x/x+12=1/x-1 ? #NoQuestionsAsk...! fo real yall be tryin my life with these can you solve it .... Naww i want the answer thats that ! if yall cant handle it dnt even bother answering it !

Parth (parthkohli):

• Multiply both sides by \(x + 12\). • Multiply both sides by \(x - 1\). You get the fractions eliminated, now solve the linear equation.

OpenStudy (anonymous):

ok sounds good now i get x^2-2x-12=0 right ?

OpenStudy (anonymous):

GO ahead..

Parth (parthkohli):

Let me see. \( \color{Black}{\Rightarrow x = \Large {x + 12 \over x + 1}}\) \( \color{Black}{\Rightarrow x^2 + x = x + 12}\) \( \color{Black}{\Rightarrow x^2 + x - x - 12 = 0}\) \( \color{Black}{\Rightarrow x^2 - 12 = 0}\)

OpenStudy (anonymous):

yes ? and ?

Parth (parthkohli):

Then you solve for \(x\).

OpenStudy (anonymous):

There is x - 1.. You are right @kliz21

OpenStudy (anonymous):

I said you @ParthKohli don't wear it on eyes...

Parth (parthkohli):

Sorry.

OpenStudy (anonymous):

my equations tell me that its ok

OpenStudy (anonymous):

Do you know how to use quadratic formula ?? @kliz21

Parth (parthkohli):

\( \color{Black}{\Rightarrow x^2 - 2x - 12 = 0}\) \[ \color{Black}{\Rightarrow x = {-(-2) \pm \sqrt{(-2)^2 - 4(1)(-12)} \over 2(1)} }\]

OpenStudy (anonymous):

lolll alex Get Outta Hea ! im trynna be nice fa once in my life !

OpenStudy (anonymous):

lol word jaja ok

OpenStudy (anonymous):

ok spiderman guy i think i Got it now and to the waterin eyse dnt i say NO QUESTIONS ASK !

Parth (parthkohli):

Good job!